A 20.0 kg child plays on a swing having support ropes that are 2.30 m long. A friend pulls her back until the ropes are 39.0 degrees from the vertical and releases her from rest.

What is the potential energy for the child just as she is released, compared with the potential energy at the bottom of the swing?

How fast will she be moving at the bottom of the swing?

How much work does the tension in the ropes do as the child swings from the initial position to the bottom?
Answer #1

a)

vertical height is given by

h = L ( 1 - cos x)

h = 2.3* ( 1 - cos 39)

h = 0.51256 m

potential energy at top

pE = mg h = 20* 9.8 * 0.51256

pE = 100.4626 J

=====

b)

using 3rd equation of motion

v^2 = 2 g h

v = 3.17 m/s

=====

c)

workdone by tension

W = T d cos 90 = 0 J

=====

Comment before rate in case any doubt, will reply for sure.. goodluck